Consider the following numerical example using the Solow growth model. Suppose that


F (K, N) = zK1/2N1/2


Furthermore, assume that 5% of the capital is lost each period due to depreciation, the population grows by 1% each period, the consumer in this economy saves 20% of his income and the total factor productivity is z = 2. The unit period is one year.


1. Find the steady state per-capita quantity of capital (k*), production (y*) and consumption (c*). [5 pts]


2. Find the steady state quantity of capital per worker that maximize consumption per worker in this model. [4 pts]


3. Derive the golden rule steady state per-capita consumption (c**), production (y**) and saving (s**). [6 pts]

1
Expert's answer
2021-12-15T07:32:01-0500

(1)

"F (K, N) = zK^{1\/2}N^{1\/2}"

Given

"\delta" =5%

"n" =1%

"s" =20%

"z" =2, "\alpha"="\frac{1}{2}"

We know,

From steady state condition:

"k^*=" steady state capital.

"=(\frac{sz}{n+\delta})^\frac{1}{1-\alpha}"

"=(\frac{0.2\times 2}{0.01+0.05})^\frac{1}{1-\frac{1}{2}}"

"=44.44"

"\therefore F(.)=z(k^*)^\alpha"

"=2\times(44.44)^\frac{1}{2}"

"y^*=13.33"

"c^*=y-s y=(1-s)y=0.8\times 13.33"

"=10.66"

(2)

From steady state condition:

"k^*=" steady state capital.

"=(\frac{sz}{n+\delta})^\frac{1}{1-\alpha}"

"=(\frac{0.2\times 2}{0.01+0.05})^\frac{1}{1-\frac{1}{2}}"

"=44.44"

(3)

Golden rule

"k_{gr}=(\frac{\alpha}{\delta+n})^\frac{1}{1-\alpha}"

"=(\frac{\frac{1}{2}}{0.05+0.01})^\frac{1}{1-\frac{1}{2}}"

"=69.44"

"y^{**}=2(69.44)^\frac{1}{2}=16.66"

"c^{**}=16.66-(0.01+0.05)69.44"

"=12.49"

"s^{**}=y^{**}-c^{**}"

"=16.66-12.49"

"=4.17"

Savings rate

"=\frac{4.17}{16.66}\times100"%"=25.03"%